Find the equation of the tangent to at the point where , in f(x)=x^2+4x-5

standard form.


1
Expert's answer
2022-03-16T17:46:22-0400

We proceed to differentiate: f'(x)=2x+4


Then we substitute for the point x0:


"y-y_0=f'(x_0)*(x-x_0)
\\ y-y_0=(2x_0+4)*(x-x_0)
\\ y=y_0+(2x)x_0-2x_0^2+4x-4x_0
\\ y=x_0^2+4x_0-5+(2x)x_0-2x_0^2+4x-4x_0
\\ y=2x(x_0+2)-x_0^2-5=mx+b"


This leads us to the conclusion that the standard form for the tangent line is the line with slope "m=2(x_0+2)" and the intersection at "b=-(x_0^2+5)", where x0 is the point where the tangent line is placed.